OpenStudy (anonymous):

Solve for x problems. will give medal!! 3 over 4 x + 5 over 8 = 4x (x plus 16)over 3 = 3x

OpenStudy (solomonzelman):

\[\frac{3}{4x}+\frac{5}{8}=4x\] you can use one of your equations, both aren't necessary for one variable

OpenStudy (anonymous):

No, they are two different equations

OpenStudy (anonymous):

I am not sure how to solve them. They are two different ones tho. @SolomonZelman

OpenStudy (solomonzelman):

1) Everything times 4x write out what you get from this.

OpenStudy (anonymous):

its 3+2.5 when multiplied by 4

OpenStudy (anonymous):

so 3x+2.5=4x

OpenStudy (solomonzelman):

Sorry, everything times 8x

OpenStudy (anonymous):

6x+5=4x

OpenStudy (solomonzelman):

6+5x=16x^2

OpenStudy (anonymous):

how did you get the 16^2

OpenStudy (solomonzelman):

4x times 4x or (4x)^2 the square goes on x and the 4, so 16x^2

OpenStudy (solomonzelman):

\[6+5x=16x^2\]

OpenStudy (anonymous):

would it be 4?

OpenStudy (solomonzelman):

the x? No! Solve it first, don't guess.

OpenStudy (solomonzelman):

subtract 16x^2 from both sides